OpenStudy (anonymous):

HOw do you find the derivative of (x^4)+(x^2)^2

OpenStudy (anonymous):

x^4 + (x^2)^2 = x^4 + x^4 = 2x^4 the derivative is 8x^3

OpenStudy (anonymous):

x^4 + (x^2)^2 = x^4 + x^4 = 2x^4 the derivative is 8x^3